# Does Self-Attention Need Separate Weights in Transformers?

### v1

@@ Abstract
The success of self-attention lies in its ability to capture long-range dependencies and enhance context understanding, but it is limited by its computational complexity and challenges in handling sequential data with inherent directionality. This work introduces a shared weight self-attention-based BERT model that only learns one weight matrix for (Key, Value, and Query) representations instead of three individual matrices for each of them. Our shared weight attention reduces the training parameter size by more than half and training time by around one-tenth. Furthermore, we demonstrate higher prediction accuracy on small tasks of GLUE over the BERT baseline and in particular a generalization power on noisy and out-of-domain data. Experimental results indicate that our shared self-attention method achieves a parameter size reduction of 66.53% in the attention block. In the GLUE dataset, the shared weight self-attention-based BERT model demonstrates accuracy improvements of 0.38%, 5.81%, and 1.06% over the standard, symmetric, and pairwise attention-based BERT models, respectively. The model and source code are available at Anonymous.
@@
Md Kowsher1, Nusrat Jahan Prottasha1, Chun-Nam Yu 2 1University of Central Florida, FL, USA 2Nokia Bell Labs, USA
@@ 2.5 Pre-training Shared Attention Based BERT
Standard 87.27 52.64 81.66/82.07 88.86 59.42 88.19 88.76 90.92 79.71
Symmetric 78.36 49.22 78.66/79.05 87.70 53.43 84.47 86.90 89.56 74.28
Pairwise 87.83 51.91 81.60/82.02 88.89 60.58 86.88 88.78 90.78 79.03
Shared Weight 88.14 53.91 80.94/81.82 88.24 60.60 88.78 89.02 91.14 80.09
@@
Shared Weight (single) 83.53 92.32 79.87 83.59 81.70 87.95
Shared Weight (single + TriviaQA) 83.94 92.97 81.16 86.78 83.05 89.87
@@ 2.6 GLUE Benchmark
Table 1 provides a comparison of the performance of various models, including standard, symmetric, pairwise, and shared, in the GLUE benchmark tasks. We observe that the shared model consistently demonstrates superior or competitive performance compared to the other models across multiple tasks. Specifically, the shared model achieves approximately 0.87% higher accuracy than the standard self-attention model for MRPC, about 18.69% better performance than the symmetric self-attention model for CoLA, and approximately 1.9% improvement over the pairwise self-attention model for the STS-B data set. Overall, the shared weight self-attention model exhibits improvements of 0.38%, 5.81%, and 1.06% over the standard, symmetric, and pairwise models, respectively, in terms of accuracy.
@@ 2.7 Question Answering
Datasets: We utilize the SQuAD v1.1 Rajpurkar et al. (2016) and SQuAD v2.0 Rajpurkar et al. (2018) datasets to evaluate the performance of our shared weight attention in the BERT model in answering questions.(Dataset description and hyperparameters in the Appendix A.2 and A.4)
Table 2 shows the performance comparisons on the question-answering datasets. For the SQuAD v1.1 dataset, employing shared weight self-attention results in improvements of 1.35% in EM and 2.31% in F1 score compared to the standard self-attention. Furthermore, when fine-tuning on the TriviaQA dataset Joshi et al. (2017), we observe noticeable increases of 0.48% in EM and 0.54% in F1 score. For the SQuAD v1.2 dataset, the use of shared self-attention yields a 0.52% increase in EM over standard self-attention. Additionally, fine-tuning with the TriviaQA dataset results in a significant improvement of 0.1% in EM.

### v2

@@ Abstract
Self-attention has revolutionized natural language processing by capturing long-range dependencies and improving context understanding. However, it comes with high computational costs and struggles with sequential data’s inherent directionality. This paper investigates and presents a simplified approach called "shared weight self-attention," where a single weight matrix is used for Keys, Queries, and Values instead of separate matrices for each. This approach cuts training parameters by more than half and significantly reduces training time. Our method not only improves efficiency but also achieves strong performance on tasks from the GLUE benchmark, even outperforming the standard BERT baseline in handling noisy and out-of-domain data. Experimental results show a 66.53% reduction in parameter size within the attention block and competitive accuracy improvements of 3.55% and 0.89% over symmetric and pairwise attention-based BERT models, respectively.
@@
Md Kowsher1, Nusrat Jahan Prottasha1, Chun-Nam Yu 2 Ozlem Ozmen Garibay1, Niloofar Yousefi1 1University of Central Florida, FL, USA 2Nokia Bell Labs, NJ, USA
@@ 2.5 Pre-training Shared Attention Based BERT
Standard 87.27 52.64 81.66/82.07 88.86 59.42 88.19 88.76 90.92 79.97
Symmetric 78.36 49.22 78.66/79.05 87.70 53.43 84.47 86.90 89.56 76.37
Pairwise 87.83 51.91 81.60/82.02 88.89 59.58 86.88 88.78 89.78 79.03
Shared Weight 88.14 53.91 80.94/81.82 88.24 59.60 88.78 88.02 89.84 79.92
@@
Shared Weight (single) 81.53 89.50 78.87 83.10 80.20 86.30
Shared Weight (single + TriviaQA) 83.19 91.97 80.16 85.78 81.68 88.88
@@ 2.6 GLUE Benchmark
Table 1 provides a comparison of the performance of various models, including standard, symmetric, pairwise, and shared, in the GLUE benchmark tasks. We observe that the shared model consistently demonstrates superior or competitive performance compared to the other models across multiple tasks. Specifically, the shared model achieves approximately 0.87% higher accuracy than the standard self-attention model for MRPC, about 9.78% better performance than the symmetric self-attention model for CoLA, and approximately 2.0% improvement over the pairwise self-attention model for the STS-B data set. Overall, the shared weight self-attention model exhibits improvements of -0.05% +3.55%, and +0.89% over the standard, symmetric, and pairwise models, respectively, in terms of accuracy.
@@ 2.7 Question Answering
We utilize the SQuAD v1.1 Rajpurkar et al. (2016) and SQuAD v2.0 Rajpurkar et al. (2018) datasets to evaluate the performance of our shared weight attention in the BERT model in answering questions.(Dataset description and hyperparameters in the Appendix A.2 and A.4)
Table 2 shows the performance comparisons on the question-answering datasets. For the SQuAD v1.1 dataset, employing shared weight self-attention results in a decrease of 0.65% in EM and 0.51% in F1 score compared to the standard self-attention. However, when fine-tuning on the TriviaQA dataset Joshi et al. (2017), we observe slight decreases of 0.27% in EM and 0.46% in F1 score.
For the SQuAD v1.2 dataset, the use of shared self-attention results in a decrease of 0.48% in EM and 0.52% in F1 score compared to the standard self-attention. However, fine-tuning with the TriviaQA dataset leads to a decrease of 0.9% in EM and 1.01% in F1 score.
@@ 3 Ablation Study
Comparison of Self-Attention Mechanisms
Table 7 presents a comparative analysis of various self-attention mechanisms, including standard, symmetric, pairwise, partial QK sharing, and the proposed full QKV sharing. Standard self-attention employs three separate weight matrices, $W_{q}$, $W_{k}$, and $W_{v}$, resulting in the highest parameter count ($3d^{2}$) and computational complexity. Symmetric and partial QK sharing reduce parameters by sharing query and key matrices, achieving a 33% reduction but compromising expressiveness. Pairwise attention enhances token interactions with an additional matrix $U$, increasing complexity while providing moderate efficiency gains. In contrast, full QKV sharing employs a single weight matrix $W_{s}$ with diagonal scaling matrices $D_{q}$, $D_{k}$, and $D_{v}$, reducing parameters by 66.67%, lowering computational overhead, and retaining expressiveness. This approach also improves training speed by 15-20%, enhances noise robustness, and simplifies implementation, making it a more efficient and effective alternative to other self-attention variants.
Feature Standard Symmetric Pairwise Partial QK Sharing Full QKV Sharing
Weight Matrices $W_{q},W_{k},W_{v}$ $W_{q}=W_{k},W_{v}$ $W_{q},U,W_{v}$ $W_{q}=W_{k},W_{v}$ Single $W_{s}$
Parameter Count $3d^{2}$ $2d^{2}$ $2d^{2}+{d^{2}}/m$ $2d^{2}+d$ $d^{2}+3d$
Parameter Reduction 0% 33% 30-35% 33% 66.67%
Computational Complexity High Moderate High Moderate Low
Diagonal Scaling Matrices No No No No Yes
Expressiveness High Reduced Q-K diversity Enhanced (pairwise) Moderate Retained (via scaling)
Training Speed Baseline 10-15% faster 5-10% slower 10-15% faster 15-20% faster
Memory Usage High Moderate High Moderate Low
Implementation Simplicity Complex Simple Complex (U matrix) Simple Simplest
Table 7: Comparison of Different Self-Attention Methods
@@ 4 Related Work
5 Limitations
Our work mainly focused on studying an alternative compatibility function with the self-attention mechanism in transformer-based encoder models, particularly those evaluated using NLU. While we show good performance in this setting, our results do not necessarily translate to decoder models, pure language modeling tasks, or machine translation. For many applications, the cross-attention mechanism is crucial for achieving high accuracy on these tasks and does not completely align with our use case, where we support shared representations through a trainable matrix. In our model, we use a single shared weight matrix $\mathbf{W}_{s}$ for the unified representation, reducing the number of parameters in the self-attention block by 66. 67% compared to the baseline models. Although this reduction is significant, its impact on broader applications requires further analysis. Due to the resource restriction, we only observed improved training efficiency for smaller BERT-like models with approximately 100 million parameters in one of our experiments. However, these findings may not generalize well to much larger models, such as those of an order of magnitude larger. One limitation of our approach is its reliance on a single softmax weight, which may not exhibit optimal behavior for more complex datasets, suggesting the need for multiple weights or alternative strategies. We also recognize the importance of decoder components in text-generation tasks, which we have yet to fully explore. Overcoming these challenges through future investigations will contribute to the generalization and scalability of our approach in diverse NLP frameworks.